OpenStudy (anonymous):

Someone help me , asap ! Ted practices two types of swimming styles for a total of 50 minutes every day. He spends 20 minutes more practicing the breast stroke than the butterfly stroke. Part A: Write a pair of linear equations to show the relationship between the number of minutes Ted practices the butterfly stroke every day (x) and the number of minutes he practices the breast stroke every day (y). Part B: How much time does Ted spend on practicing the breast stroke every day? Part C: Is it possible for Ted to have spent 45 minutes practicing the butterfly stroke every day?

OpenStudy (anonymous):

Someone please help me with this one, now. It's urgently needed.

OpenStudy (anonymous):

y=x+20 x+y=50 Butterfly = 15 Breast Stroke = 35 Part C = No

OpenStudy (anonymous):

Let butterfly stroke = x Breaststroke = y 20 more minutes on breast than butterfly means Breast is Butterfly +20. or Y = x+20 The total time he practices 50 is both styles added together. Time on butterfly = x PLUS time on Breaststroke = y or y+x=50

OpenStudy (anonymous):